|x-1| + |x-2| + |x-8| + |x-9| =f(x) find range of function

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
If
x>= 9
Then
x-1 + x-2 + x-8 +x-9 = f(x)
4x -20 = f(x)
f(x) = [16, infinity)
Now
if 8
x-1 + x-2 + x-8 + 9-x = f(x)
2x-2 = f(x)
f(x) = [14, 16]
Now if 2
14 = f(x)
Now if 1
18 - 2x = f(x)
f(x) = [16, 14]
If x
20 - 4x = f(x)
f(x) = (infinity, 16]